A section of the Uncanoonuc Mountains, called the West Side Trail, is located in New Hampshire. The trail has a vertical change of about 268 feet vertically for every 1000 feet it moves along Horizontally. What is the slope of the trail?
slope=rate of change= (change in y)/ (change in x)= (vertical change)/(horizontal change)
slope= 268 feet/1000ft=268/1000=67/250

Found 2 solutions by checkley77, eperette:
Answer by checkley77(12844)

(0,0)(1000,268)
slope=(y2-y1)/(x2-x1)=(268-0)/(1,000-0)=268/1000=.268 is the slope.

Answer by eperette(173)